OpenStudy (anonymous):

If the side length of a square can be represented by 2x + 10 and its area is 256 square units, find the value of x.

OpenStudy (anonymous):

(2x+10)^2=256 4x^2 + 40x + 100 = 256 can you solve from there?

OpenStudy (anonymous):

I got this far but I am having problems what to do from here....having a mental block because I know it is something simple I am missing.

OpenStudy (anonymous):

i hate this problem... um subtract 256 from both sides next but I don't know what after that. ^^;

OpenStudy (anonymous):

me too. you have to subtract 100 from both sides but then I can't remember how to get 4x^ +40x=156 ( I think) ....solving for x

OpenStudy (anonymous):

I got this far but I am having problems what to do from here....having a mental block because I know it is something simple I am missing.